the perimeter of a rectangle is 72 mm. the height is three more than two times the base. what is the area
Answer by stanbon(75887)

the perimeter of a rectangle is 72 mm.
the height is three more than two times the base.
---
Let the base be "x".
Then the height is "2x+3"
----
Perimeter = 2(base + height)
-----
72 = 2(x + 2x+3)
3x+3 = 36
3x = 33
x = 11 mm (base)
2x+3 = 2*11+3 = 25 mm (height)
===========================
what is the area ?
Area = base *height
Area = 11*25 = 275 mm^2
===========================
Cheers,
Stan H.